The probabilities of a machine manufacturing 0, 1, 2, 3, 4, or 5 defective parts in one day are 0.75, 0.17, 0.04, 0.025, 0.01, and 0.005, respectively. Find the mean of the probability distribution. 



1
Expert's answer
2021-02-03T15:20:27-0500

Solution:


First of all, let us restate the formula of the mean:


"E(\xi) = \Sigma_{i=1}^n p(x_i)x_i"


Mean of the given probability distribution:


"E(\xi) = p(0)*0 + p(1)*1 + p(2)*2 + p(3)*3 + p(4) *4 +p(5)*5 ="

"= 0.17 * 1 + 0.04*2+0.025*3+0.01*4+0.005*5 ="

"=0.17+0.08+0.075+0.04+0.025=0.17+0.12+0.1="

"=0.39"


Answer:


"E(\xi) = 0.39"